SX AO Reconstructor Measurement

Summary

Details

03:00 Start up WFS on SX and DX

04:09 Restore board 8 Feb 16 W_internal
  • See light on CCD47 adn CCD39

04:10 No light on DX WFS (still)

4:49 Light on Pyramid from RR, CLosed loop, offload
  • Hex posiiton -2.93, 2.242, 4.76. 138.9, -166.4

05:00 Check cube rot postion
  • start with 182.181
  • finish 182.171

05:11 Closed loop 400 mode
  • get instability with ho2 = 0.25

05:17 Rotate instrument
  • 280, -60.76, -32.63
  • 270, -60.48, -36.81
  • 260, -58.14, -42.02
  • 250, -54.97, -45.36
  • 240, -49.88, -49.85
  • 230, -44.45, -53.54
  • 220, -39.29, -55.85
  • 210, -33.76, -58.71
  • 200, -25.56, -60.40
  • 190, -20.93, -60.06
  • 180, -11.21, -58.99
  • 170, -5.10, -57.17
  • 160, 1.33, -52.19
  • 150, 9.15, -48.09
  • 140, 14.37, -39.86
  • 130, 17.48, -31.78
  • 120, 22.01, -25.06
  • 110, 22.34, -16.04
  • 100, 20.81, -7.66

05:35 Increase gains at rot=100, see same instability in ho2 = 0.25

05:40 Rotate back to 283
  • loose star on CCD47
  • have to go out of focus and change rx, ry
  • get coma
  • close loop, T/T. let offload, add ho1
  • -2.93, 2.242, 4.77, 140.178, -170.4
  • bayside -73.5, -51.2

Bin 1, pupil 20120909-001205

06:05 Disturb
  • 10 modes
  • KL_v18
  • 4 fames
  • 45 pp
  • 180512
  • wrong gain file

06:05 Disturb
  • 10 modes
  • KL_v18
  • 4 frames
  • optmized_amano_3lambda
  • 45 pp
  • 180722

06:06 Acquire
  • 12 lam/d = 6.7V
  • open loop
  • 4 iter
  • 180930

06:10 Analyze
  • failed

06:12 Acquire
  • 12 lam/d = 6.7V
  • open loop
  • 4 iter
  • 181203

06:13 Analyze
  • 181346

06:15 Acuire
  • 12 lam/d = 6.7V
  • closed loop
  • 4 iter
  • 181644

06:18 Analyze
  • 181828

06:21 Disturb
  • 66 modes
  • KL_v18
  • 4 frames
  • optmized_amano_3lambda
  • 7 pp
  • 182114

06:23 Acquire
  • 3 lam/d = 2.28V
  • closed loop
  • 4 iter
  • 182433

06:27 Analyze
  • 182646

06:27 Load default.sav. Gets rid of "dot" at 4:30 in flat

06:28 Acquire
  • use 66 mode recon 182646
  • closed loop
  • 66 mode
  • 182936

06:31 Analyze
  • 183106

06:33 Disturb
  • 153 modes
  • KL_v18
  • 4 frames
  • optmized_amano_3lambda
  • 3 pp
  • 183316

06:34 Acquire
  • use 66 mode recon 182106
  • closed loop
  • 153 mode
  • 183535

06:38 Analyze
  • 183719

06:39 Acquire
  • use 153 mode recon 182719
  • closed loop
  • 153 mode
  • 183951

Actuator 1, 2, 6, 14 removed from position

Actuator 2, 14 have no commands

06:44 Analyze
  • 184236

06:45 Distrub
  • 4 iter
  • 1 pp
  • 400 modes
  • 184539

06:49 Acquire
  • use 153 mode recon
  • 4 iter
  • 184825

06:51 Analyze
  • 185106

07:05 Acquire
  • use 400 mode recon 185106
  • 400 modes
  • 48 iter
  • 190920

07:25 Analyze
  • 192325

07:33 CL with 192325
  • Optical Loop data 193316
  • over driven 193405

07:39 CL
  • recon 192325
  • disturbance with 0.8" seeing and Gopt Astig 20160209_164809
  • 193852
  • ratio on Gopt GUI is 0.69
  • re-normalize gains 0.66 to optical gain and ratio goes to 1.
  • Apply to Gopt and apply to Slop\e Computer
  • opt loop data 194219
  • apply 100 nm Z5 astig to DM. Can see in Pupil images
  • 194320

Bin 3

07:47 Disturb
  • 10 modes
  • 4 frames
  • 45 pp
  • opamized_amano_6l
  • 194729

07:48 Acquire
  • open loop
  • 600 Hz
  • 6 lam/D
  • 4 iter
  • 194856

07:50 Analyze
  • 195101

07:53 Acquire
  • use 10 mode recon 195101
  • closed loop
  • 4 iter
  • 195401

07:55 Analyze
  • 195528

07:56 Test and Optical Gain to 1.0
  • Reset Gopt "set button

08:10 Test 20160309_195514
  • 201318
  • 1.2" seeing

08:16 Disturb
  • 4 iter
  • 7 pp
  • bin 3
  • 66 modes
  • 201709

08:19 Acquire
  • 4 iter
  • 66 modes
  • 201842

08:20 Analyze
  • 202010

08:23 Acquire
  • bin 3
  • 48 iter
  • 202303

08:37 Analyze
  • 203720

08:42 Test Bin 3
  • 1.2" seeing
  • 204209 Optical loop data

Check pupil alignment

08:45 Both horizonal and vertical alignment looks the same as two days ago...good

Check AO patrol field

09:07 Test SX (
  • 0,0 = -77.5, -49.
  • 77, 0 = -0.5, -49.
  • -34, 0 = -1115, -49.1
  • 0, -34 = -77.5, -85.1
  • 0, 49 = -0.1, -77.5

09:34 Moveable telescope
  • preset to BS9107
  • ACE_TRACK mode
  • RS=0
  • Left 0.27, 0.04
  • Right 1.55, 2.13
  • failed, CCD47 not powered
  • Powered it

09:36 Preset, GS=0
  • Left 0.27, 0.04 = -77.2, -49
  • Right 1.55, 2.13 = -75.0, -48.6

09:43 Preset GS=1
  • left 23.66, 2.91 = -53.83, -46.19
  • right 24.94, 5.0 = -98, -45

09:50 preset TS=4, RS=1 (4 is east
  • left OK
  • right fail

09:50 preset TS=1, RS=4 (4 is west
  • left fail
  • right OK

Test RETROREF focal station values

10:33 TT = 0 volts. Apply Loop Params then ok at 2.28

10:39 Now Hexapod at -3.76, 2.525, 4.73, 82, -344

-- %USERSIG{DougMiller - 2016-02-20}%
Topic revision: r3 - 20 Feb 2016, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback